A rectangle has a perimeter of 80 cm. If its width is x, express its length and its area in terms of x. What is the maximum area of the rectangle?
Answer by stanbon(75887)

y^2-4y-45 = 0
-------
I'm not sure what method your teacher used.
Your problem is factorable as follows.
This is sometimes called inverse-FOIL.
-----
y^2-9y+5y-45 = 0
------
Factor the 1st 2 and the last 2 terms separately:
y(y-9)+5(y-9) = 0
------
Factor again:
(y-9)(y+5) = 0
---
Solve:
y = 9 or y = -5
======================
Cheers,
Stan H.
============